the mean of 5 number is 20.if 1 number is excluded mean of the remaning number becomes23

Answer by fcabanski(1391)

Sum / # of numbers = mean.


sum / 5 = 20


sum = 5*20 = 100


sum = mean * number of numbers.
For the four remaining numbers, the sum is 4 * 23 = 92. The number removed was 100-92 = 8.